### v2

# Improving Intrusion Detection with Domain-Invariant Representation Learning in Latent Space

Abstract
Domain generalization focuses on leveraging knowledge from multiple related domains with ample training data and labels to enhance inference on unseen in-distribution (IN) and out-of-distribution (OOD) domains. In our study, we introduce a two-phase representation learning technique using multi-task learning. This approach aims to cultivate a latent space from features spanning multiple domains, encompassing both native and cross-domains, to amplify generalization to IN and OOD territories. Additionally, we attempt to disentangle the latent space by minimizing the mutual information between the prior and latent space, effectively de-correlating spurious feature correlations. Collectively, the joint optimization will facilitate domain-invariant feature learning. We assess the model’s efficacy across multiple cybersecurity datasets, using standard classification metrics on both unseen IN and OOD sets, and juxtapose the results with contemporary domain generalization methods.
Keywords:
Domain generalization Representation learning Multi-task learning
1 Introduction
Domain generalization seeks to learn from multiple source domains and subsequently assess the model on an unseen target domain [18, 19]. A domain typically constitutes a probability distribution where training samples originate. These samples stem from a source domain, while a separate target domain is kept for testing. Deep neural networks excel when tested on data from the same distribution as their training data. Within the few-shot learning (FSL) framework, models aim to adapt to out-of-distribution (OOD) data with minimal labeled samples [10, 9, 8, 7, 6]. Meta-learning endeavors to train models across varied tasks, enabling them to grasp new tasks with few samples [11]. A pivotal challenge arises from the domain shift between training and test distributions, compromising generalization [12]. Many recent studies train with both in-distribution and OOD data, though this can infuse bias. Dataset bias emerges when a model trained on one dataset falters on test data from a distinct distribution [21]. This is commonly known as domain adaptation where information about the target domain or its classes is accessible during the training process[22]. Some approaches augment training data with perturbations mimicking domain shifts. In domain generalization, the underlying assumption is that there’s no foreknowledge of the target distribution. The challenge lies in designing a model that, once trained on single or multiple datasets, can classify data in unfamiliar domains without further training. We aspire to devise features that bridge multiple related domains without directly engaging them during the training process. Latent space models like Autoencoders function by encoding inputs into a lower-dimensional latent space and subsequently decoding them to get an estimate of the original input. A significant inspiration for our work is denoising autoencoders, which aim to reconstruct original data from noise-altered input, thereby bolstering encoder robustness. Previous research has proposed methods such as minimizing divergence between data source marginal distributions or adversarial feature alignment using maximum mean discrepancy [21]. We focus on learning a disentangled latent representation by consistently minimizing mutual information between prior and latent spaces. This action indirectly tries to maximize the mutual independence between input and latent spaces. Our model adopts a supervised autoencoder (SAE) approach [15], where both classification and reconstruction processes occur in the learned latent space simultaneously. However, our multi-domain latent space is curated with samples spanning multiple other domains(cross-domains) with the mutual invariance loss trying to remove the spurious correlation of the original high dimensional data. Recently, the work [37] has highlighted the problem of how spurious correlation can hurt generalization performance when the task correlation changes among tasks and solves the problem with multi-task causal representation learning with task-specific invariant regularization. In our case, we consider the scenario where we do not know the features with spurious correlation beforehand and therefore, directly minimize the MI content between the prior and the latent space which indirectly will aim to remove the spurious correlation in the latent space and create a disentangled latent space. This is based on the intuition that most probably the spurious correlation of the high-dimensional input space will be transferred to the latent space if there is no proper regularization to minimize it. Multi-task learning will facilitate learning representations from multiple diverse datasets and therefore jointly optimizing it with a de-correlation penalty between input and latent space will enable learning a disentangled or invariant latent space representation which will help us achieve generalization to unseen domains. In summary, our contributions can be highlighted as:
  • We introduce a two-stage representation learning approach for domain generalization in the latent space. In the first phase, we craft a multi-domain latent space by involving the data from multiple source and cross-domains. Simultaneously, we try to cultivate a disentangled or invariant latent space by applying a mutual information(MI) regularization which tries to minimize the MI between the high-dimensional prior space and the latent space. The regularization acts as a de-correlation penalty by removing spurious correlation that exists in different attack datasets in the original high-dimensional input space. These two components jointly try to enforce domain invariant learning in the latent space.
  • Our methodology jointly optimizes the classification loss, the multi-domain reconstruction loss, and the mutual invariance regularization in the latent space. This joint optimization seeks to boost classification outcomes, particularly considering both IN and OOD classes.
  • We perform a wide range of experiments on cybersecurity datasets. Our results showcase overall significant improvement in generalization performance for all the source, cross-domain, and OOD classes when evaluated with standard classification metrics.
2 Related Work
Domain generalization techniques can be grouped into the following primary categories: domain invariant representation learning, meta-learning,latent dimension regularization, and metric learning.
1) Domain Invariant Representation Learning: This method aims to identify domain invariant representations that can be extended to unseen domains. The crux of these strategies, as seen in works such as [35], is to filter out domain-specific insights while maintaining cross-domain information. Notable studies employing autoencoders, such as [1], amalgamate multiple domains during training, augmented by data enhancement techniques, to extract domain-invariant characteristics. These features then demonstrate superior generalization to out-of-distribution data. Another study, Maximum Mean Discrepancy Adversarial Autoencoder (MMD-AAE) [3], in the context of few-shot learning, emphasizes aligning varied domain distributions to a generic prior distribution while engaging in adversarial feature learning. An innovative approach is suggested in [4], where a domain-centric masking technique is applied to learn both domain-specific and domain-invariant features. This will facilitate efficient source domain classification and sufficient generalization to target domains. In [15], a noise-enhanced supervised autoencoder is presented. This autoencoder undertakes a dual task of reconstructing inputs and classifying both the inputs and the reconstructed entities by feeding them back as input to the model. Leveraging the intra-class correlation (ICC) metric, they demonstrate the superior discriminative and generalizing capabilities of the learned feature embeddings. Moreover, authors in [26] propose domain generalization through domain-invariant representation that uniformly distributes across multiple source domains. Their approach employs moment alignment of distributions and enforces feature disentanglement via an entropy loss.
2) Meta-learning: This approach employs learning from a number of related tasks for domain generalization, as observed in works such as [11, 25, 31]. The study in [27] introduces a technique to discern a domain interdependent projection leading to a latent space. This space minimizes biases in the data while preserving the inherent relationship across multiple domains. Model Agnostic Meta-Learning (MAML) has also been extended to latent dimension settings by performing the gradient-based adaptation in the low dimensional space instead of the higher dimensional space of model parameters [30]. Zero-shot learning [32] aims at learning models from seen classes and inferring on samples whose categories were unseen during the training process.
3) Information Bottleneck Principle: In contrast to the aforementioned methodologies, our strategy propels direct disentanglement or decorrelation between multiple training domains. Decorrelating spurious feature correlations in order to isolate signals from background noise has been successfully used in the physics and statistics community before. An information-theoretic perspective on variance-invariance-covariance has been provided here [29] in the context of self-supervised learning which helps to achieve generalization guarantees for downstream supervised learning tasks. Adversarial learning-based domain adaptation methods are prone to negative transfer which hurts the generalization performance [28]. In our method, class-invariant features are directly learned in the latent dimension by minimizing mutual information between latent and prior space. Utilizing kernel-based metrics, we non-parametrically estimate mutual information between latent and prior space in order to achieve this.
4) Metric Learning: Metric learning aims to learn a representation function that can map higher-dimensional data to a latent embedded space. The distance in embedded space should preserve the similarity between the similar classes and the dissimilar classes are pushed away from each other. The authors [36] propose mixing target labels with training samples to improve the quality of representations or embeddings for classification purposes. In deep metric learning, we try to learn non-linear mapping from input space to low-dimensional latent space [33]. Various mix-up techniques have been recently proposed to improve the quality of representation learned by adopting some interpolation techniques between pairs of input samples and their labels [34]. On the other hand, we follow a multi-task representation learning approach that aims to learn a robust feature representation in the latent space by leveraging multiple sources and cross-domain feature information.
3 Problem Formulation
In our domain generalization problem, we consider $M$ training tasks comprising of source and cross-domain data consisting of $N$ samples, $\mathcal{S}_{train}=\{X^{i},Y^{i}|i=1,2,3...,M\}$ where, ${X}^{i},{Y}^{i}=\{\left(x^{i}_{j},y^{i}_{j}\right)\}_{j=1}^{N_{i}}$. Let $I=\{I_{1},I_{2},...,I_{k}\}$ represent the power set of these labels, where $I_{q}\subseteq\{0,1,...,K\}$. For example, if $I_{1}$ is a source domain, where, $I_{1}=\{0,1,2,3,4\}$, then, we add multiple cross-domain data $I_{2},I_{3},I_{4}$, where, $I_{2}=\{5,6,7\}$, $I_{3}=\{8,9,10\}$, and $I_{4}=\{11,12,13\}$. Let’s assume that in this context, the cardinality of each $I_{q}$, denoted as $|I_{q}|$, is constrained such that $|I_{q}|\leq 5$. The extension to multiple domains necessitates the definition of a multi-task learning objective over all the M source and cross domains which can be given as
$\begin{split}\mathcal{L}_{rec}\left({S}_{train};\theta,\phi\right)=\sum_{i=1}^{M}\Big{\|}f_{\theta}^{(i)}\left(g_{\phi}^{(i)}\left(X^{i}\right)\right)-X^{i}\Big{\|}^{2}_{2}\\ \end{split}$ (1)
In this expression, $g_{\phi}^{(i)}$ and $f_{\theta}^{(i)}$ denote the encoder and decoder functions respectively for each of the $M$ sources and cross domains, $X^{i}$ is the input training data from a particular domain. The total reconstruction loss is the sum of reconstruction losses over the source and cross-domain latent space reconstructions. The cross-domain data aims to introduce more diversity in the latent space representation which in turn can help to achieve better generalization performance.
3.1 Mutual Invariance Regularization
In order to create a domain-invariant latent space, we use a nonparametric way to estimate the mutual information between the prior and the latent space directly from the samples.
3.1.1 Using Renyi’s Entropy and Joint Entropy to Calculate Mutual Information Between Inp and Latent Space
In information theory, the dependence measure or the total correlation quantifies how much the feature variables either in latent space or input space, deviate from the statistical independence in each dimension and is expressed as the Kullback Leibler(KL) divergence between the joint probability distribution and the marginal distribution of the feature variables[b16]. In our case, we enforce the de-correlation between the high-dimensional input space and low-dimensional latent space representing multiple source and cross-domain data by minimizing the mutual information between the input and the latent space as a regularization technique. This will help in creating a disentangled latent space that is devoid of spurious correlations of the high-dimensional input space.
For a given data matrix $X^{l}$, where $l$ represents a batch size, the matrix-based Renyi’s $\alpha$ entropy for a normalized positive semi-definite matrix ${X}$ of size $n\times n$, where $n$ represents the number of samples in each batch, can be estimated in the input space as
$\begin{split}\hat{H_{2}}(X^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(X^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(X^{l})^{\alpha}\right),\end{split}$ (2)
where $\left(X^{1}\right)_{i,j}=\mathcal{K}_{1}\left(x^{1}_{i},x^{1}_{j}\right)$, $\left(X^{2}\right)_{i,j}=\mathcal{K}_{2}\left(x^{2}_{i},x^{2}_{j}\right)$ ,…, $\left(X^{l}\right)_{i,j}=\mathcal{K}_{l}\left(x^{l}_{i},x^{l}_{j}\right)$ and $\lambda_{i}({X})$ denotes the $i^{th}$ eigenvalue of $X$, $x^{l}$ represents the $l_{th}$ batch.
Similarly, Renyi’s quadratic entropy of the encoded latent space $Z$ is estimated as
$\begin{split}\hat{H_{2}}(Z^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(Z^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(Z^{l})^{\alpha}\right),\end{split}$ (3)
The argument in equation (3) is called the information potential. In the above section, we use the matrix-based second-order Renyi’s entropy $\left(\alpha=2\right)$ [16] to evaluate the entropy of the latent and prior space. This can be expressed in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the Hilbert space. We use this measure to estimate the entropy of the original data space. Now, we can write the matrix-based $\alpha$ order joint entropy between latent space $Z$ and prior space $X$ as
$\hat{H_{2}}\left(X^{l},Z^{l}\right)=H_{2}\left(\frac{X^{l}\circ Z^{l}}{tr\left(X^{l}\circ Z^{l}\right)}\right),$ (4)
where $\circ$ represents the Hadamard product of matrices. Based on the above definitions, we calculate the mutual information between latent and prior space with the help of the matrix-based normalized Renyi’s entropy of the latent space and the joint entropy between the prior space and latent space directly from the samples.
3.1.2 The Mutual Information Divergence
We use the non-parametric method of mutual information divergence in order to estimate the mutual information between the latent and prior space. Minimizing the mutual information indirectly results in de-correlating the feature correlation that exists in the original input space and the latent space which helps in improving the generalization performance. The mutual information during each batch of the training can be estimated as
$\begin{split}\hat{MI}(X^{l};Z^{l})&=\hat{H_{2}}\left(X^{l}\right)+\hat{H_{2}}\left(Z^{l}\right)-\hat{H_{2}}\left(X^{l},Z^{l}\right),\end{split}$ (5)
where $\hat{H_{2}}(X,Z)$ is the second-order joint entropy between latent and prior space. Minimizing this divergence as a regularization technique in the final loss objective will aid in preserving useful disentangled information in the latent space during each iteration of the training process.
3.2 The Multi-Task Learning Objective
In the unsupervised learning setting using latent space models, it is assumed that the label or the class information is not readily available. However, to achieve OOD generalization, we leverage the label information of the multiple source and cross-domain encoded data in the latent space during the training process.
In our approach, we do a joint optimization of the classification and the reconstruction loss using the latent space encoding. The total loss calculated over all the $N$ samples during the training process can be given as
$\begin{split}\mathcal{L}\left(\mathcal{S}_{train},Z;{\theta},{\phi}\right)=\min_{\phi,\theta}\frac{1}{N}\sum_{i=1}^{N}\Bigr{\{}\mathcal{L}^{\left(i\right)}_{ce}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ +\beta\mathcal{L}^{(i)}_{MI}\left(X^{i};Z^{i}_{\theta,\phi}\right)+\lambda\mathcal{L}_{rec}\left(X^{i};\theta,\phi\right)\Bigl{\}},\end{split}$ (6)
where, $\mathcal{L}_{ce}$ is the cross-entropy loss calculated on the latent space encoding considering the binary classification problem, given as,
$\mathcal{L}_{ce}^{(i)}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ =-\left(y^{i}\log\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\\ +\left(1-y^{i}\right)\log\left(1-\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\right)\right)$
$\mathcal{L}_{MI}$ is the disentanglement or decorrelation loss between the latent space $Z$ and the input space data $X$ in the form of mutual information as given in eq: 5, $\mathcal{S}_{y}$ is the softmax function applied on the encoded data $g_{\phi}(x)$, $\mathcal{L}_{rec}$ is the reconstruction loss. Here, we augment the cross-entropy loss with the mutual information regularization between the latent and prior space. The regularization aims to penalize spurious information content between the encoding and the prior space and the strength of the penalty is determined by the hyper-parameter $\beta$. $\beta$ controls the trade-off between decorrelation and classification. During the joint optimization, we try to balance the reconstruction weight and the mutual information regularization.
4 Experiments
In this section, we demonstrate the performance of our proposed model on benchmark cybersecurity datasets.
4.1 Dataset
  • CSE-CIC-IDS2018 This is a publicly available cybersecurity dataset that is made available by the Canadian Cybersecurity Institute (CIC). We combine the data from different days which has different classes of anomalies to get a multi-domain dataset. While testing, we consider the data from a different day which consists of other distinct classes of normal and anomalies as our OOD data.
  • NSL-KDD This is also a publicly available benchmark cybersecurity dataset made available by CIC. It has a total of 43 different features of internet traffic flow. The experimental results of our analysis are reported in our github repository.
4.2 Baselines
We consider the following models related to latent space multi-task representation learning and few-shot learning as baselines in order to compare and evaluate the performance of our proposed model.
  • Deep-CNN This is a supervised CNN model trained on all the features. The number of convolution, pooling, and fully connected layers are adjusted in order to achieve the best performance on all the classes. We perform an ensembling approach and calculate the average accuracy after running multiple iterations.
  • Deep Multi-task AE [1] In this latent dimension encoder-decoder-based model the reconstruction error is optimized over multiple domains in an unsupervised fashion. We consider multiple sources and cross-domain data and train them jointly with the label information in latent space. This is basically our model without the de-correlation regularization.
  • MMD-AE [3, 2] This paper uses the Minimum Mean Discrepancy (MMD) measure as regularization for domain generalization between multiple cross-domain data. We use it as a few-shot learning method where the cross-domain data are added to improve the OOD generalization.
  • Noise Enhanced Supervised AE (NSAE) [15]This model jointly predicts the labels of inputs and reconstructs the inputs considers reconstructed samples as noisy inputs and feeds them back to the model. In this model, it requires to do an extra step of fine-tuning by further training the reconstructed data using a supervised classifier.
  • MAML [11] For this baseline, we consider data samples from different domains to form the final classification task and evaluate the performance on unseen domain data (OOD). We consider the samples from the cross-domain classes to fine-tune the generalization process using the bi-level optimization algorithm.
4.3 Training Strategy
In the training phase, we jointly optimize multiple tasks by classifying multiple source and cross-domain samples from different datasets in the latent space. In our experiments with CIC-CSE-IDS data, our source domain data consists of the attack classes like GOLDEN-EYE, SOLARIS, and the BENIGN class, whereas our cross-domain samples are taken from attack classes like INFILTRATION, BRUTFORCE, and BOTNET. During the testing phase, we consider the OOD attack classes like DDOS-HOIC, RARE, and SlowHTTPTest along with an OOD BENIGN class to evaluate our domain generalization performance. During training, we introduce the cross-domain data in various proportions ranging from 5%-35% of the total source domain attack samples. This allows the model to learn a diverse latent space representation by reconstructing different source and cross-domain data while also simultaneously de-correlating spurious correlation between input and latent space by enforcing the mutual-information penalty. We use a learning rate of $0.005$ and train the model for $300$ epochs for all the training scenarios. We use a stacked encoder-decoder architecture which gradually projects the high-dimensional input space to low-dimensional latent space step-wise. The code is implemented in Pytorch and standard optimization libraries such as Adam are used during the training process.  [^1] https://anonymous.4open.science/r/MTRAE-40FA/
Algorithm 1 The Multi-task Latent Space Regularized Encoder-Decoder Model (MTLS-RED)
Input: Source domain data ${X_{1},X_{2},...,X_{m}}$, ${X}\in\mathcal{R}^{d}$, $\forall{m}\in\{1,2,3,..\}$,
Add cross-domain data ${X_{1}^{\prime},X_{2}^{\prime},...,X_{m}^{\prime}}$, ${X^{\prime}}\in\mathcal{R}^{d}$, $\forall{m}\in\{4,5,6,..,\}$,
Let the source and cross-domain labels be $\{y_{i}^{m}\}_{i=1}^{n}$, $\forall{m}\in\{1,2,3,4,5,6,...,M\}$
Initialize the encoder(E) and decoder(D) weights: $\mathbf{W}_{\phi}\in\mathcal{R}^{d_{x}\times d_{z}}$, $\mathbf{W}_{\theta}\in\mathcal{R}^{d_{z}\times d_{x}}$ , learning rate $\alpha$
while not end of epochs do:
     for batch = 1 to total batches do:
         Sample mini-batch data $\{X_{i}\}_{1}^{l}\in\mathbf{R}^{d}$, l is batch-size
         Compute the RBF kernels of prior space $K_{x_{l}}$ and latent space $K_{z_{l}}$
         Calculate mutual information between input space $X_{l}$ and latent space $Z_{l}$ using matrix-based Renyi’s entropy ${MI\left(X_{l}||Z_{l}\right)}$.
         Perform a forward pass on encoder ${E\left(X_{{\phi}_{i}}\right)}$
         Calculate total batch loss $\mathcal{L}_{l}=\mathcal{L}_{ce}\left({X_{i},y_{i}}\right)+\mathcal{L}_{rec}\left({D\left(E\left(X_{i}\right)\right),X_{i}}\right)+\lambda\mathcal{L_{MI}}\left({X_{i}}||{Z_{i}}\right)$
         Update $\mathbf{W}_{\phi}$ and $\mathbf{W}_{\theta}$
         $\mathbf{W}_{\phi_{t+1}}\leftarrow\mathbf{W}_{\phi_{t}}-\alpha_{1}\nabla_{\phi}\mathcal{L}\left(\theta,\phi\right)$
         $\mathbf{W}_{\theta_{t+1}}\leftarrow\mathbf{W}_{\theta_{t}}-\alpha_{2}\nabla_{\theta}\mathcal{L}\left(\theta,\phi\right)$
     end for
end while
Output: Trained MTLS-RED algorithm
Figure 1: Fig shows MI Gain($\beta$) variation with classification metrics on SOURCE and CROSS DOMAIN data (SOLARIS, GOLDENEYE, BENIGN,BRUT-FORCE, INFIL, BOTNET) and OOD DATA(RARE, HOIC, SLOW-HTTPS, BENIGN). We keep reconstruction loss weightage fixed. Fig shows that maximum accuracy is obtained when MI is between 1 and 3 for all the cases.
4.4 Sensitivity Analysis
In our joint optimization objective, we need to balance weightage to the hyperparameters $\beta$ and $\lambda$ in order to achieve good generalization performance to all the classes. We find that allocating a higher weightage to the entropy loss, $\beta=1.5$ to $\beta=3$, and keeping the reconstruction loss weightage at $\lambda=0.5$ to $\lambda=0.6$ yields the best performance of the model considering all the source, cross-domain, and OOD datasets. Fig 1 shows the maximum accuracy for most of the SOURCE, CROSS, and OOD domains achieved at this range. Another important hyper-parameter is the kernel bandwidth $\sigma$ used to estimate the matrix-based Renyi’s quadratic entropy in the latent space and prior space. We experiment with various kernel widths and report the best results in Table 1 and Table 2. We consider the batch size in between $200-300$ samples during each step of the gradient descent training as MI is easier to estimate when the sample size is relatively high.
ACCURACY-IN ACCURACY-OOD
Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
MAML $93.50$ $99.70$ $69.14$ $73.50$ $30.31$ $86.03$ $54.24$ $34.58$
Deep-CNN $\mathbf{100.00}$ ${\mathbf{100.00}}$ 99.0599.05{99.05}99.05 50.9550.95{50.95}50.95 $36.90$ $89.16$ $58.56$ $71.34$
D-MTAE $99.50$ $99.60$ $\mathbf{99.40}$ $99.50$ $\mathbf{61.30}$ 27.3827.38{27.38}27.38 $32.20$ $61.00$
MMD-AAE $99.34$ $99.70$ 99.0499.04{99.04}99.04 $99.11$ $48.05$ $28.00$ 02.0002.00{02.00}02.00 $46.25$
NSAE $\mathbf{100.00}$ $\mathbf{100.00}$ $55.07$ 50.0050.00{50.00}50.00 48.7648.76{48.76}48.76 $\mathbf{100.00}$ $50.56$ 61.3961.39{61.39}61.39
MTLS-RED 88.8388.83{88.83}88.83 86.0886.08{86.08}86.08 82.4182.41{82.41}82.41 $\mathbf{96.12}$ 28.0928.09{28.09}28.09 80.9980.99{80.99}80.99 $\mathbf{73.93}$ $\mathbf{73.23}$
Table 1: Table 1 shows the accuracy of different SOURCE domains comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET. The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN. In this case, we add 10% of source domain attack samples as cross-domain data.
ACCURACY-IN ACCURACY-OOD
Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
MAML $97.35$ $97.40$ $54.84$ $87.51$ $46.39$ $70.45$ $34.78$ $70.45$
Deep-CNN $\mathbf{100.00}$ $\mathbf{100.00}$ $\mathbf{99.94}$ 51.4151.41{51.41}51.41 $36.92$ $61.65$ $49.52$ $77.81$
D-MTAE $61.30$ $61.30$ $32.20$ $61.00$ $72.60$ $48.40$ $69.80$ $\mathbf{83.32}$
MMD-AAE $96.61$ $99.46$ 99.2199.21{99.21}99.21 $78.23$ $24.84$ $72.08$ 47.1347.13{47.13}47.13 $46.25$
NSAE 99.9899.98{99.98}99.98 $\mathbf{100.00}$ $34.85$ 49.4249.42{49.42}49.42 48.4048.40{48.40}48.40 $\mathbf{82.70}$ $\mathbf{99.25}$ 51.5951.59{51.59}51.59
MTLS-RED $79.46$ $79.33$ $79.20$ $\mathbf{79.12}$ $\mathbf{78.99}$ 79.3379.33{79.33}79.33 $79.25$ 79.6479.64{79.64}79.64
Table 2: Table 2 shows the accuracy of the SOURCE data comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET. The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN. In this case, we add 35% of source domain attack samples as cross-domain data.
4.5 Ablation Study
The ablation study is performed keeping the baseline models as reference. In order to study the effectiveness of our proposed model, we perform the following analysis
  • We consider the multiple source and cross-domain reconstructions and use the encoded space to classify SOURCE, CROSS-DOMAIN, and OOD data. Our setup is different from [1] where the authors adopt a two-stage approach where the model is trained first, and then the trained encoder is used for classification purposes. On the other hand, our strategy is to jointly train all the components together so that the joint optimization with the de-correlation penalty can improve the generalization performance. The joint objective promotes discriminative power while simultaneously reducing the dependency on spurious dataset-specific correlation.
  • In our experiments we increase the cross-domain data availability gradually from 5%-35% in the training data in order to understand the impact of decorrelation penalty in achieving domain generalization while reconstructing data from multiple domains. In Table 1, we observe that when the cross-domain attack classes constitute a small percentage $10\%$ of the source domain data, most of the baseline models overfit the source domain classes and do not perform well on the cross-domain and test OOD classes. The OOD classes are never involved in the training process. Whereas, our proposed model gives an overall good performance on all the source, cross-domain, and OOD classes except the RARE class. In Table 2, we report the results of the experiments where we gradually increase the cross-domain data to $35\%$. In this case, with our proposed model, the best accuracy generalizes well to all the source, cross-domain, and OOD classes although there is a decrease in performance with some of the source and cross-domain classes but better performance can be seen with the OOD classes especially. In a nutshell, we want to point out that by adjusting the kernel bandwidth and the sample size, we can get a reliable estimate of the mutual information penalty between prior and latent space which can help achieve better generalization to completely unseen classes that were never used in the training process.
5 Conclusion
In this paper, we attempt to study an important problem of domain generalization where a model trained on several sources and cross-domains can generalize better on OOD classes(’zero-day’ attacks in the case of cyber-security). We showcase that learning feature representation from multiple diverse domains in the latent space with proper regularization can achieve better OOD generalization. We formulate a joint objective that tries to classify data in a multi-domain latent space while simultaneously enforcing a de-correlation penalty in the form of mutual information between input and latent space to create a domain-invariant latent space. Through detailed experiments, we showcase the performance of the model on benchmark cybersecurity datasets with the help of standard classification metrics. In the future, we would like to study domain generalization in time series and other sequential data using latent space regularization.

### v3

@@
-Improving Intrusion Detection with Domain-Invariant Representation Learning in Latent Space
+Out-of-Distribution Detection with Domain-Invariant Representations in Multi-domain Latent Space
 Abstract
-Domain generalization focuses on leveraging knowledge from multiple related domains with ample training data and labels to enhance inference on unseen in-distribution (IN) and out-of-distribution (OOD) domains.
-In our study, we introduce a two-phase representation learning technique using multi-task learning.
-This approach aims to cultivate a latent space from features spanning multiple domains, encompassing both native and cross-domains, to amplify generalization to IN and OOD territories.
-Additionally, we attempt to disentangle the latent space by minimizing the mutual information between the prior and latent space, effectively de-correlating spurious feature correlations.
+Domain generalization focuses on leveraging the knowledge from the training data of multiple related domains to enhance inference on unseen in-distribution (IN) and out-of-distribution (OOD) domains.
+In our study, we introduce a multi-task representation learning technique that leverages the information of multiple related domains to improve the detection of classes from unseen domains.
+Our method aims to cultivate a latent space from data spanning multiple domains, encompassing both source and cross-domains, to amplify generalization to OOD domains.
+Additionally, we attempt to disentangle the latent space by minimizing the mutual information between the input and the latent space, effectively de-correlating spurious correlations among the samples of a specific domain.
 Collectively, the joint optimization will facilitate domain-invariant feature learning.
-We assess the model’s efficacy across multiple cybersecurity datasets, using standard classification metrics on both unseen IN and OOD sets, and juxtapose the results with contemporary domain generalization methods.
+We assess the model’s efficacy across multiple cybersecurity datasets, using standard classification metrics on both unseen IN and OOD sets, and validate the results with contemporary domain generalization methods.
 Keywords:
@@
 1 Introduction
-Domain generalization seeks to learn from multiple source domains and subsequently assess the model on an unseen target domain [18, 19].
-A domain typically constitutes a probability distribution where training samples originate.
+Domain generalization seeks to learn from multiple related source domains and subsequently evaluate the model on an unseen target domain [18, 19].
+A domain typically constitutes a particular probability distribution from which the training samples are drawn.
 These samples stem from a source domain, while a separate target domain is kept for testing.
 Deep neural networks excel when tested on data from the same distribution as their training data.
-Within the few-shot learning (FSL) framework, models aim to adapt to out-of-distribution (OOD) data with minimal labeled samples [10, 9, 8, 7, 6].
+Within the few-shot learning (FSL) framework, models aim to adapt to out-of-distribution (OOD) data with minimal labeled samples [10, 8, 7, 6].
 Meta-learning endeavors to train models across varied tasks, enabling them to grasp new tasks with few samples [11].
@@
 Many recent studies train with both in-distribution and OOD data, though this can infuse bias.
-Dataset bias emerges when a model trained on one dataset falters on test data from a distinct distribution [21].
-This is commonly known as domain adaptation where information about the target domain or its classes is accessible during the training process[22].
-Some approaches augment training data with perturbations mimicking domain shifts.
+In the case of domain adaptation, it is considered that the information about the target domain or its classes is accessible during the training process [22].
+Some approaches augment training data with perturbations mimicking the target domain.
 In domain generalization, the underlying assumption is that there’s no foreknowledge of the target distribution.
 The challenge lies in designing a model that, once trained on single or multiple datasets, can classify data in unfamiliar domains without further training.
-We aspire to devise features that bridge multiple related domains without directly engaging them during the training process.
-Latent space models like Autoencoders function by encoding inputs into a lower-dimensional latent space and subsequently decoding them to get an estimate of the original input.
-A significant inspiration for our work is denoising autoencoders, which aim to reconstruct original data from noise-altered input, thereby bolstering encoder robustness.
-Previous research has proposed methods such as minimizing divergence between data source marginal distributions or adversarial feature alignment using maximum mean discrepancy [21].
-We focus on learning a disentangled latent representation by consistently minimizing mutual information between prior and latent spaces.
-This action indirectly tries to maximize the mutual independence between input and latent spaces.
-Our model adopts a supervised autoencoder (SAE) approach [15], where both classification and reconstruction processes occur in the learned latent space simultaneously.
-However, our multi-domain latent space is curated with samples spanning multiple other domains(cross-domains) with the mutual invariance loss trying to remove the spurious correlation of the original high dimensional data.
-Recently, the work [37] has highlighted the problem of how spurious correlation can hurt generalization performance when the task correlation changes among tasks and solves the problem with multi-task causal representation learning with task-specific invariant regularization.
-In our case, we consider the scenario where we do not know the features with spurious correlation beforehand and therefore, directly minimize the MI content between the prior and the latent space which indirectly will aim to remove the spurious correlation in the latent space and create a disentangled latent space.
-This is based on the intuition that most probably the spurious correlation of the high-dimensional input space will be transferred to the latent space if there is no proper regularization to minimize it.
-Multi-task learning will facilitate learning representations from multiple diverse datasets and therefore jointly optimizing it with a de-correlation penalty between input and latent space will enable learning a disentangled or invariant latent space representation which will help us achieve generalization to unseen domains.
+We aspire to devise domain invariant features that bridge multiple related domains without directly engaging them during the training process.
+Previous research has proposed methods such as adversarial feature alignment using maximum mean discrepancy [21] to achieve domain generalization.
+In our case, we focus on learning a disentangled latent representation by consistently minimizing mutual information between prior and latent spaces from multiple related domains.
+However, our multi-domain latent space is curated with samples spanning multiple other domains(cross-domains) with the mutual invariance penalty trying to remove the spurious correlation present in individual domains.
+Recently, the work [37] has highlighted the problem of how spurious correlation can hurt generalization performance when the task correlation changes among the related tasks and solves the problem with multi-task causal representation learning with task-specific invariant regularization.
+In our case, we consider the scenario where the samples from individual domains have spurious feature correlation and therefore minimize the mutual information content between the input and the latent space which indirectly will aim to remove the spurious correlation in the individual domains.
+Otherwise, this is done by measuring the entropy or the uncertainty of the data on the eigenspectrum of the gram matrix of the input and latent space kernels.
+Multi-task learning facilitates learning representations from multiple diverse datasets and therefore jointly optimizing it with a de-correlation penalty to achieve generalization to unseen domains.
 In summary, our contributions can be highlighted as:
-  • We introduce a two-stage representation learning approach for domain generalization in the latent space.
-In the first phase, we craft a multi-domain latent space by involving the data from multiple source and cross-domains.
-Simultaneously, we try to cultivate a disentangled or invariant latent space by applying a mutual information(MI) regularization which tries to minimize the MI between the high-dimensional prior space and the latent space.
-The regularization acts as a de-correlation penalty by removing spurious correlation that exists in different attack datasets in the original high-dimensional input space.
-These two components jointly try to enforce domain invariant learning in the latent space.
+  • We consider a scenario where we know that each domain has its own spuriously correlated features which hurts the generalization performance of the model when tested on OOD domains.
+To address this, we propose a multi-task representation learning approach that adapts a multi-domain latent space by incorporating data from multiple sources and cross-domains.
+Simultaneously, we try to cultivate a disentangled or invariant latent space by applying a regularization penalty that minimizes the mutual information(MI) between the input and the latent space.
   • Our methodology jointly optimizes the classification loss, the multi-domain reconstruction loss, and the mutual invariance regularization in the latent space.
-This joint optimization seeks to boost classification outcomes, particularly considering both IN and OOD classes.
-  • We perform a wide range of experiments on cybersecurity datasets.
-Our results showcase overall significant improvement in generalization performance for all the source, cross-domain, and OOD classes when evaluated with standard classification metrics.
+The regularization, calculated in the kernel space of the input and latent space, acts as a de-correlation penalty, or otherwise, we do not want the spurious correlation of the higher dimensional features to be projected to the latent space.
+  • We show that cross-domain data when added in a principled way, can improve generalization performance on the IN and OOD classes.
+We perform a wide range of experiments on cybersecurity datasets to showcase overall improvement in generalization performance across IN, OOD, and cross-domains classes.
 2 Related Work
@@
 Their approach employs moment alignment of distributions and enforces feature disentanglement via an entropy loss.
+Recently, the DIFEX [9] paper utilizes a knowledge distillation framework to capture the high-level Fourier phase as internally-invariant characteristics, while simultaneously learning cross-domain correlation alignment to extract mutually-invariant features.
 2) Meta-learning: This approach employs learning from a number of related tasks for domain generalization, as observed in works such as [11, 25, 31].
@@
 Zero-shot learning [32] aims at learning models from seen classes and inferring on samples whose categories were unseen during the training process.
-3) Information Bottleneck Principle: In contrast to the aforementioned methodologies, our strategy propels direct disentanglement or decorrelation between multiple training domains.
-Decorrelating spurious feature correlations in order to isolate signals from background noise has been successfully used in the physics and statistics community before.
+3) Information Bottleneck Principle and Metric Learning: In contrast to the aforementioned methodologies, our strategy propels direct disentanglement or decorrelation between multiple training domains.
 An information-theoretic perspective on variance-invariance-covariance has been provided here [29] in the context of self-supervised learning which helps to achieve generalization guarantees for downstream supervised learning tasks.
 Adversarial learning-based domain adaptation methods are prone to negative transfer which hurts the generalization performance [28].
-In our method, class-invariant features are directly learned in the latent dimension by minimizing mutual information between latent and prior space.
-Utilizing kernel-based metrics, we non-parametrically estimate mutual information between latent and prior space in order to achieve this.
-4) Metric Learning: Metric learning aims to learn a representation function that can map higher-dimensional data to a latent embedded space.
+In our method, class-invariant features are indirectly learned in the latent dimension by minimizing mutual information between latent and prior space.
+Metric learning aims to learn a representation function that can map higher-dimensional data to a latent embedded space.
 The distance in embedded space should preserve the similarity between the similar classes and the dissimilar classes are pushed away from each other.
@@
 For example, if $I_{1}$ is a source domain, where, $I_{1}=\{0,1,2,3,4\}$, then, we add multiple cross-domain data $I_{2},I_{3},I_{4}$, where, $I_{2}=\{5,6,7\}$, $I_{3}=\{8,9,10\}$, and $I_{4}=\{11,12,13\}$.
-Let’s assume that in this context, the cardinality of each $I_{q}$, denoted as $|I_{q}|$, is constrained such that $|I_{q}|\leq 5$.
 The extension to multiple domains necessitates the definition of a multi-task learning objective over all the M source and cross domains which can be given as
@@
 3.1 Mutual Invariance Regularization
-In order to create a domain-invariant latent space, we use a nonparametric way to estimate the mutual information between the prior and the latent space directly from the samples.
-3.1.1 Using Renyi’s Entropy and Joint Entropy to Calculate Mutual Information Between Inp and Latent Space
-In information theory, the dependence measure or the total correlation quantifies how much the feature variables either in latent space or input space, deviate from the statistical independence in each dimension and is expressed as the Kullback Leibler(KL) divergence between the joint probability distribution and the marginal distribution of the feature variables[b16].
-In our case, we enforce the de-correlation between the high-dimensional input space and low-dimensional latent space representing multiple source and cross-domain data by minimizing the mutual information between the input and the latent space as a regularization technique.
-This will help in creating a disentangled latent space that is devoid of spurious correlations of the high-dimensional input space.
-For a given data matrix $X^{l}$, where $l$ represents a batch size, the matrix-based Renyi’s $\alpha$ entropy for a normalized positive semi-definite matrix ${X}$ of size $n\times n$, where $n$ represents the number of samples in each batch, can be estimated in the input space as
-$\begin{split}\hat{H_{2}}(X^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(X^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(X^{l})^{\alpha}\right),\end{split}$ (2)
-where $\left(X^{1}\right)_{i,j}=\mathcal{K}_{1}\left(x^{1}_{i},x^{1}_{j}\right)$, $\left(X^{2}\right)_{i,j}=\mathcal{K}_{2}\left(x^{2}_{i},x^{2}_{j}\right)$ ,…, $\left(X^{l}\right)_{i,j}=\mathcal{K}_{l}\left(x^{l}_{i},x^{l}_{j}\right)$ and $\lambda_{i}({X})$ denotes the $i^{th}$ eigenvalue of $X$, $x^{l}$ represents the $l_{th}$ batch.
-Similarly, Renyi’s quadratic entropy of the encoded latent space $Z$ is estimated as
-$\begin{split}\hat{H_{2}}(Z^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(Z^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(Z^{l})^{\alpha}\right),\end{split}$ (3)
+In order to learn a domain-invariant latent space, we use a kernel-based mutual information minimization between the prior and the latent space as a regularization technique.
+3.1.1 Kernel-based Renyi’s Entropy and Joint Entropy to measure Mutual Information between Input and Latent Space
+In information theory, the dependence measure or the total correlation between the feature variables is measured as the statistical independence in each dimension and is expressed as the Kullback Leibler(KL) divergence between the joint probability distribution and the marginal distribution of the features [16].
+In our case, we enforce the de-correlation between the input and the latent kernel space comprising multiple source and cross-domains by minimizing the mutual information between them as a regularization penalty.
+This will try to introduce a disentangled latent space that is devoid of the spurious correlations in the individual domains in their high-dimensional input space.
+The matrix-based Renyi’s second-order entropy [16] of a normalized positive definite(NPD) matrix $\mathcal{K}_{x}$, of size $l\times l$ in the input space, where $l$ represents a batch size, can be estimated as
+$\begin{split}\hat{H_{2}}(\mathcal{K}_{x})=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{l}\lambda_{i}(\mathcal{K}_{x})^{\alpha}\right),\end{split}$ (2)
+where the Gram matrix $\mathcal{K}_{x}$ is obtained by evaluating the positive definite (PSD) kernel on all $l$ pairs of samples in a batch of training, that is, $\left(\mathcal{K}_{x}\right)_{i,j}=\mathcal{K}_{x}\left(x_{i},x_{j}\right)$ and $\lambda_{i}({X})$ denotes the $i^{th}$ eigenvalue of the input kernel matrix $\mathcal{K}_{x}$ of the $l_{th}$ batch, $\mathcal{K}_{i,j}=\frac{1}{l}\frac{\mathcal{K}_{i,j}}{\sqrt{\mathcal{K}_{i,i}\mathcal{K}_{j,j}}}$ Here, $\alpha=2$ considering Renyi’s second-order entropy.
+Similarly, Renyi’s quadratic entropy of the latent space kernel $\mathcal{K}_{Z}$ of size $l\times l$ is estimated as
+$\begin{split}\hat{H_{2}}(\mathcal{K}_{z})=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{l}\lambda_{i}(\mathcal{K}_{z})^{\alpha}\right),\end{split}$ (3)
 The argument in equation (3) is called the information potential.
-In the above section, we use the matrix-based second-order Renyi’s entropy $\left(\alpha=2\right)$ [16] to evaluate the entropy of the latent and prior space.
-This can be expressed in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the Hilbert space.
-We use this measure to estimate the entropy of the original data space.
-Now, we can write the matrix-based $\alpha$ order joint entropy between latent space $Z$ and prior space $X$ as
-$\hat{H_{2}}\left(X^{l},Z^{l}\right)=H_{2}\left(\frac{X^{l}\circ Z^{l}}{tr\left(X^{l}\circ Z^{l}\right)}\right),$ (4)
-where $\circ$ represents the Hadamard product of matrices.
-Based on the above definitions, we calculate the mutual information between latent and prior space with the help of the matrix-based normalized Renyi’s entropy of the latent space and the joint entropy between the prior space and latent space directly from the samples.
+In the above section, we use the matrix-based second-order Renyi’s entropy $\left(\alpha=2\right)$ [16] to evaluate the entropy or the uncertainty of the latent and the input space in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the Hilbert space.
+Now, we can estimate the matrix-based second-order joint entropy between latent space kernel $Z$ and the input space kernel $X$ as
+$\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)=H_{2}\left(\frac{\mathcal{K}_{x}\circ\mathcal{K}_{Z}}{tr\left(\mathcal{K}_{X}\circ\mathcal{K}_{z}\right)}\right),$ (4)
+where $\circ$ represents the Hadamard product.
+Based on the above definitions, we calculate the joint entropy of the latent and the input space with the help of the matrix-based normalized Renyi’s entropy of the latent space and the input space kernels.
+The joint entropy is used to derive the mutual information between the input and the latent space.
 3.1.2 The Mutual Information Divergence
-We use the non-parametric method of mutual information divergence in order to estimate the mutual information between the latent and prior space.
-Minimizing the mutual information indirectly results in de-correlating the feature correlation that exists in the original input space and the latent space which helps in improving the generalization performance.
+We use the matrix-based mutual information divergence to estimate the mutual information between the latent and input space kernels.
+Minimizing the mutual information indirectly results in de-correlating the feature correlation that exists in the original input space which helps in improving the generalization performance.
 The mutual information during each batch of the training can be estimated as
-$\begin{split}\hat{MI}(X^{l};Z^{l})&=\hat{H_{2}}\left(X^{l}\right)+\hat{H_{2}}\left(Z^{l}\right)-\hat{H_{2}}\left(X^{l},Z^{l}\right),\end{split}$ (5)
-where $\hat{H_{2}}(X,Z)$ is the second-order joint entropy between latent and prior space.
-Minimizing this divergence as a regularization technique in the final loss objective will aid in preserving useful disentangled information in the latent space during each iteration of the training process.
+$\begin{split}\hat{MI}(\mathcal{K}_{X};\mathcal{K}_{Z})&=\hat{H_{2}}\left(\mathcal{K}_{X}\right)+\hat{H_{2}}\left(\mathcal{K}_{Z}\right)-\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right),\end{split}$ (5)
+where $\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right),$ is the second-order joint entropy between the latent and the input kernel space.
+Minimizing this divergence as a regularization penalty in the final loss objective will aid in preserving useful disentangled information in the latent space during each iteration of the training process.
 3.2 The Multi-Task Learning Objective
-In the unsupervised learning setting using latent space models, it is assumed that the label or the class information is not readily available.
-However, to achieve OOD generalization, we leverage the label information of the multiple source and cross-domain encoded data in the latent space during the training process.
-In our approach, we do a joint optimization of the classification and the reconstruction loss using the latent space encoding.
-The total loss calculated over all the $N$ samples during the training process can be given as
-$\begin{split}\mathcal{L}\left(\mathcal{S}_{train},Z;{\theta},{\phi}\right)=\min_{\phi,\theta}\frac{1}{N}\sum_{i=1}^{N}\Bigr{\{}\mathcal{L}^{\left(i\right)}_{ce}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ +\beta\mathcal{L}^{(i)}_{MI}\left(X^{i};Z^{i}_{\theta,\phi}\right)+\lambda\mathcal{L}_{rec}\left(X^{i};\theta,\phi\right)\Bigl{\}},\end{split}$ (6)
+In our latent space multi-task learning approach, we leverage the label information of the multiple source and cross-domain encoded data in the latent space during the training process.
+In our approach, we do a joint optimization of the classification and the reconstruction loss along with the MI regularization in the latent space.
+The total loss calculated over all the $N$ batches, each consisting of $l$ samples, during the training process can be given as
+$\begin{split}\mathcal{L}\left(\mathcal{S}_{train},Z;{\phi},{\theta}\right)=\min_{\phi,\theta}\frac{1}{N}\sum_{l=1}^{N}\Bigr{\{}\mathcal{L}_{ce}\left(g_{\phi}\left(X^{l}\right),y^{l}\right)\\ +\beta\mathcal{L}_{MI}\left(X^{l};Z^{l}\right)+\lambda\mathcal{L}_{rec}\left(X^{l};\phi,\theta\right)\Bigl{\}},\end{split}$ (6)
 where, $\mathcal{L}_{ce}$ is the cross-entropy loss calculated on the latent space encoding considering the binary classification problem, given as,
-$\mathcal{L}_{ce}^{(i)}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ =-\left(y^{i}\log\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\\ +\left(1-y^{i}\right)\log\left(1-\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\right)\right)$
-$\mathcal{L}_{MI}$ is the disentanglement or decorrelation loss between the latent space $Z$ and the input space data $X$ in the form of mutual information as given in eq: 5, $\mathcal{S}_{y}$ is the softmax function applied on the encoded data $g_{\phi}(x)$, $\mathcal{L}_{rec}$ is the reconstruction loss.
+$\mathcal{L}_{ce}\left(g_{\phi}\left(X\right),y\right)\\ =-\left(y\log\left(\mathcal{S}_{y}\left(g_{\phi}\left(X\right)\right)\right)\\ +\left(1-y\right)\log\left(1-\left(\mathcal{S}_{y}\left(g_{\phi}\left(X\right)\right)\right)\right)\right)$
+$\mathcal{L}_{MI}$ is the disentanglement or de-correlation loss between the latent space and the input space expressed in the form of mutual information divergence measured in their kernel space, given in eq: 5, $\mathcal{S}_{y}$ is the softmax function applied on the encoded data $g_{\phi}(x)$, $\mathcal{L}_{rec}$ is the reconstruction loss, $\phi$, $\theta$ are the encoder and decoder parameters.
 Here, we augment the cross-entropy loss with the mutual information regularization between the latent and prior space.
-The regularization aims to penalize spurious information content between the encoding and the prior space and the strength of the penalty is determined by the hyper-parameter $\beta$.
+The regularization aims to penalize spurious information content between the latent encoding and the input space, the strength of which is determined by the hyper-parameter $\beta$.
 $\beta$ controls the trade-off between decorrelation and classification.
@@
   • CSE-CIC-IDS2018 This is a publicly available cybersecurity dataset that is made available by the Canadian Cybersecurity Institute (CIC).
-We combine the data from different days which has different classes of anomalies to get a multi-domain dataset.
-While testing, we consider the data from a different day which consists of other distinct classes of normal and anomalies as our OOD data.
-  • NSL-KDD This is also a publicly available benchmark cybersecurity dataset made available by CIC.
-It has a total of 43 different features of internet traffic flow.
-The experimental results of our analysis are reported in our github repository.
+It consists of 7 major kinds of intrusion datasets.
+We use SOLARIS, GOLDENEYE as source domain data, INFILTRATION, BOTNET as cross-domain data, and RARE, SLOWHTTPS, HOIC as OOD data, along with a OOD BENIGN class from a different day.
+  • CICIoT 2023 This is a state-of-the-art dataset for profiling, behavioral analysis, and vulnerability testing of different IoT devices with different protocols from the network traffic, consisting of 7 major attack classes.
+We use BENIGN, DoS, and DDoS as source data, RECON, as cross-domain data and WEB, MIRAI as OOD test data.
+  • CICIoMT 2024 This is a benchmark dataset to enable the development and evaluation of Internet of Medical Things (IoMT) security solutions.
+The attacks are categorized into five classes.
+We use BENIGN, DDoS, DoS as source-domain, RECON, and SPOOFING as cross-domain, and MQTT as OOD data.
+ [^1] CIC- Canadian Institute of Cybersecurity
 4.2 Baselines
-We consider the following models related to latent space multi-task representation learning and few-shot learning as baselines in order to compare and evaluate the performance of our proposed model.
-  • Deep-CNN This is a supervised CNN model trained on all the features.
-The number of convolution, pooling, and fully connected layers are adjusted in order to achieve the best performance on all the classes.
-We perform an ensembling approach and calculate the average accuracy after running multiple iterations.
-  • Deep Multi-task AE [1] In this latent dimension encoder-decoder-based model the reconstruction error is optimized over multiple domains in an unsupervised fashion.
+We consider the following models related to latent space multi-task representation learning and few-shot learning as baselines to compare and evaluate the performance of our proposed model.
+More details about the baselines can be found in the supplementary material.
+  • Correlation Alignment for Deep Domain Adaptation (CORAL) [39] This paper adopts an unsupervised domain adaptation technique that tries to align the covariance of the source and the target distributions.
+We extend CORAL to latent space and align the source and cross-domain data correlation to improve generalization on OOD data.
+  • Deep Multi-task Autoencoder (MTAE) [1] In this latent dimension encoder-decoder-based model the reconstruction error is optimized over multiple domains in a supervised fashion.
 We consider multiple sources and cross-domain data and train them jointly with the label information in latent space.
 This is basically our model without the de-correlation regularization.
-  • MMD-AE [3, 2] This paper uses the Minimum Mean Discrepancy (MMD) measure as regularization for domain generalization between multiple cross-domain data.
+  • Minimum Mean Discrepancy-Autoencoder(MMD-AE)[3, 2] This paper uses the MMD measure as regularization for domain generalization between multiple cross-domain data.
 We use it as a few-shot learning method where the cross-domain data are added to improve the OOD generalization.
-  • Noise Enhanced Supervised AE (NSAE) [15]This model jointly predicts the labels of inputs and reconstructs the inputs considers reconstructed samples as noisy inputs and feeds them back to the model.
-In this model, it requires to do an extra step of fine-tuning by further training the reconstructed data using a supervised classifier.
-  • MAML [11] For this baseline, we consider data samples from different domains to form the final classification task and evaluate the performance on unseen domain data (OOD).
-We consider the samples from the cross-domain classes to fine-tune the generalization process using the bi-level optimization algorithm.
+  • Noise Enhanced Supervised Autoencoder (NSAE) [15]This model jointly predicts the labels of inputs and reconstructs the inputs considers reconstructed samples as noisy inputs and feeds them back to the model.
+In this model, requires doing an extra step of fine-tuning by further training the reconstructed data using a supervised classifier.
 4.3 Training Strategy
-In the training phase, we jointly optimize multiple tasks by classifying multiple source and cross-domain samples from different datasets in the latent space.
-In our experiments with CIC-CSE-IDS data, our source domain data consists of the attack classes like GOLDEN-EYE, SOLARIS, and the BENIGN class, whereas our cross-domain samples are taken from attack classes like INFILTRATION, BRUTFORCE, and BOTNET.
-During the testing phase, we consider the OOD attack classes like DDOS-HOIC, RARE, and SlowHTTPTest along with an OOD BENIGN class to evaluate our domain generalization performance.
-During training, we introduce the cross-domain data in various proportions ranging from 5%-35% of the total source domain attack samples.
-This allows the model to learn a diverse latent space representation by reconstructing different source and cross-domain data while also simultaneously de-correlating spurious correlation between input and latent space by enforcing the mutual-information penalty.
-We use a learning rate of $0.005$ and train the model for $300$ epochs for all the training scenarios.
-We use a stacked encoder-decoder architecture which gradually projects the high-dimensional input space to low-dimensional latent space step-wise.
-The code is implemented in Pytorch and standard optimization libraries such as Adam are used during the training process.
+In the training phase, we jointly optimize multiple tasks by considering multiple source and cross-domain data in the latent space.
+We consider a setting where the joint distribution of the data doesn’t have much overlap but the marginal distribution of the features do have an overlap.
+This scenario happens when the features exhibit unique correlation patterns between the source and cross-domain data distributions.
+For example, DOS and DDOS datasets have similar correlation patterns between the features.
+However, the the features in the other attacks such as INFIL and WEB have different correlation patterns.
+Similarly, the GOLDEN and SOLARIS datasets have a certain kind of correlation between the features that differ from other attack datasets like INFIL and BOTNET.
+Our goal is to combine the datasets with different patterns of correlation structure among its features during the training phase in order to increase the versatility of the latent space.
+For example, if two or more features of any two attack datasets exhibit distinct correlation patterns, we add one as the source domain and the other as cross-domain data.
+This will tend to introduce more variety in the input space and more generalized information in the latent space after applying the mutual information penalty.
+During training, we vary the cross-domain data in various proportions ranging from 5%-50% of the source domain samples and evaluate the results on the OOD data.
+This allows the model to learn a diverse latent space representation by reconstructing different source and cross-domain data while simultaneously de-correlating domain-specific spurious correlation information.
+Figure 1: Fig (a-e) Correlation plots of various cyber-security datasets. (a)DOS (b) DDOS (c) WEB (d) INFIL (e) HOIC
  [^2] https://anonymous.4open.science/r/MTRAE-40FA/
@@
 while not end of epochs do:
-     for batch = 1 to total batches do:
+     for batch = 1 to total batches N do:
          Sample mini-batch data $\{X_{i}\}_{1}^{l}\in\mathbf{R}^{d}$, $l$ is batch-size
-         Compute the RBF kernels of prior space $K_{x_{l}}$ and latent space $K_{z_{l}}$
-         Calculate mutual information between input space $X_{l}$ and latent space $Z_{l}$ using matrix-based Renyi’s entropy ${MI\left(X_{l}||Z_{l}\right)}$.
+         Compute the RBF kernels of input space $\mathcal{K}_{x_{l}}$ and latent space $\mathcal{K}_{z_{l}}$ of size $l\times l$
+         Calculate mutual information between input space $X_{l}$ and latent space $Z_{l}$ using matrix-based Renyi’s entropy ${MI\left(\mathcal{K}_{{X}_{l}};\mathcal{K}_{{Z}_{l}}\right)}$.
          Perform a forward pass on encoder ${E\left(X_{{\phi}_{i}}\right)}$
-         Calculate total batch loss $\mathcal{L}_{l}=\mathcal{L}_{ce}\left({X_{i},y_{i}}\right)+\mathcal{L}_{rec}\left({D\left(E\left(X_{i}\right)\right),X_{i}}\right)+\lambda\mathcal{L_{MI}}\left({X_{i}}||{Z_{i}}\right)$
+         Calculate total batch loss $\mathcal{L}_{l}=\mathcal{L}_{ce}\left({X^{l},y^{l}}\right)+\mathcal{L}_{rec}\left(X^{l},X^{{l}^{\prime}}\right)+\lambda\mathcal{L_{MI}}\left({X}^{l}||{{Z}^{l}}\right)$
          Update $\mathbf{W}_{\phi}$ and $\mathbf{W}_{\theta}$
-         $\mathbf{W}_{\phi_{t+1}}\leftarrow\mathbf{W}_{\phi_{t}}-\alpha_{1}\nabla_{\phi}\mathcal{L}\left(\theta,\phi\right)$
-         $\mathbf{W}_{\theta_{t+1}}\leftarrow\mathbf{W}_{\theta_{t}}-\alpha_{2}\nabla_{\theta}\mathcal{L}\left(\theta,\phi\right)$
+         $\mathbf{W}_{\phi_{t+1}}\leftarrow\mathbf{W}_{\phi_{t}}-\alpha_{1}\nabla_{\phi}\mathcal{L}_{l}\left(\theta,\phi\right)$
+         $\mathbf{W}_{\theta_{t+1}}\leftarrow\mathbf{W}_{\theta_{t}}-\alpha_{2}\nabla_{\theta}\mathcal{L}_{l}\left(\theta,\phi\right)$
      end for
@@
 Output: Trained MTLS-RED algorithm
-Figure 1: Fig shows MI Gain($\beta$) variation with classification metrics on SOURCE and CROSS DOMAIN data (SOLARIS, GOLDENEYE, BENIGN,BRUT-FORCE, INFIL, BOTNET) and OOD DATA(RARE, HOIC, SLOW-HTTPS, BENIGN).
-We keep reconstruction loss weightage fixed.
-Fig shows that maximum accuracy is obtained when MI is between 1 and 3 for all the cases.
-4.4 Sensitivity Analysis
-In our joint optimization objective, we need to balance weightage to the hyperparameters $\beta$ and $\lambda$ in order to achieve good generalization performance to all the classes.
-We find that allocating a higher weightage to the entropy loss, $\beta=1.5$ to $\beta=3$, and keeping the reconstruction loss weightage at $\lambda=0.5$ to $\lambda=0.6$ yields the best performance of the model considering all the source, cross-domain, and OOD datasets.
-Fig 1 shows the maximum accuracy for most of the SOURCE, CROSS, and OOD domains achieved at this range.
-Another important hyper-parameter is the kernel bandwidth $\sigma$ used to estimate the matrix-based Renyi’s quadratic entropy in the latent space and prior space.
-We experiment with various kernel widths and report the best results in Table 1 and Table 2.
-We consider the batch size in between $200-300$ samples during each step of the gradient descent training as MI is easier to estimate when the sample size is relatively high.
-ACCURACY-IN ACCURACY-OOD
-Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
-MAML $93.50$ $99.70$ $69.14$ $73.50$ $30.31$ $86.03$ $54.24$ $34.58$
-Deep-CNN $\mathbf{100.00}$ ${\mathbf{100.00}}$ 99.0599.05{99.05}99.05 50.9550.95{50.95}50.95 $36.90$ $89.16$ $58.56$ $71.34$
-D-MTAE $99.50$ $99.60$ $\mathbf{99.40}$ $99.50$ $\mathbf{61.30}$ 27.3827.38{27.38}27.38 $32.20$ $61.00$
-MMD-AAE $99.34$ $99.70$ 99.0499.04{99.04}99.04 $99.11$ $48.05$ $28.00$ 02.0002.00{02.00}02.00 $46.25$
-NSAE $\mathbf{100.00}$ $\mathbf{100.00}$ $55.07$ 50.0050.00{50.00}50.00 48.7648.76{48.76}48.76 $\mathbf{100.00}$ $50.56$ 61.3961.39{61.39}61.39
-MTLS-RED 88.8388.83{88.83}88.83 86.0886.08{86.08}86.08 82.4182.41{82.41}82.41 $\mathbf{96.12}$ 28.0928.09{28.09}28.09 80.9980.99{80.99}80.99 $\mathbf{73.93}$ $\mathbf{73.23}$
-Table 1: Table 1 shows the accuracy of different SOURCE domains comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET.
-The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN.
-In this case, we add 10% of source domain attack samples as cross-domain data.
-ACCURACY-IN ACCURACY-OOD
-Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
-MAML $97.35$ $97.40$ $54.84$ $87.51$ $46.39$ $70.45$ $34.78$ $70.45$
-Deep-CNN $\mathbf{100.00}$ $\mathbf{100.00}$ $\mathbf{99.94}$ 51.4151.41{51.41}51.41 $36.92$ $61.65$ $49.52$ $77.81$
-D-MTAE $61.30$ $61.30$ $32.20$ $61.00$ $72.60$ $48.40$ $69.80$ $\mathbf{83.32}$
-MMD-AAE $96.61$ $99.46$ 99.2199.21{99.21}99.21 $78.23$ $24.84$ $72.08$ 47.1347.13{47.13}47.13 $46.25$
-NSAE 99.9899.98{99.98}99.98 $\mathbf{100.00}$ $34.85$ 49.4249.42{49.42}49.42 48.4048.40{48.40}48.40 $\mathbf{82.70}$ $\mathbf{99.25}$ 51.5951.59{51.59}51.59
-MTLS-RED $79.46$ $79.33$ $79.20$ $\mathbf{79.12}$ $\mathbf{78.99}$ 79.3379.33{79.33}79.33 $79.25$ 79.6479.64{79.64}79.64
-Table 2: Table 2 shows the accuracy of the SOURCE data comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET.
-The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN.
-In this case, we add 35% of source domain attack samples as cross-domain data.
+4.4 Hyperparameter Sensitivity
+In the joint optimization objective, we aim to balance the weightage to the hyperparameters $\beta$ and $\lambda$ to achieve good generalization performance to all the classes.
+We find that allocating a higher weightage to the entropy loss and keeping less weightage to the reconstruction loss results in the best overall performance of the model considering different scenarios.
+The kernel bandwidth used to estimate the matrix-based Renyi’s quadratic entropy in the latent space and input space proves to be a significant hyperparameter.
+Optimum bandwidth selection is directly linked to the smoothness of the learning.
+A small bandwidth can lead to under-smoothing whereas a large bandwidth may result in over-smoothing We vary the kernel bandwidth during our experiments.
+We consider the correlated features in the case of each dataset.
+For a fair comparison between the baselines, we keep the latent space dimension similar while experimenting on the datasets with a particular baseline.
+Datasets Features learning rate kernel bw batch size Weightage Latent Dim
+CSE-CIC-IDS 79 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $79-30-15$
+CICIoT 2022 43 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $43-27-13$
+CICIoMT 2024 44 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $44-22-11$
+SOURCE DOMAIN CROSS DOMAIN OOD DOMAIN
+Model Percentage SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC HTTPS BENIGN
+DMTAE $0\%$ $\mathbf{99.97}$ 52.7052.70{52.70}52.70 04.0004.00{04.00}04.00 09.5009.50{09.50}09.50 $00.30$ $00.38$ $02.30$ 97.9897.98{97.98}97.98
+$15\%$ 99.5099.50{99.50}99.50 $\mathbf{99.70}$ $\mathbf{99.40}$ $\mathbf{99.50}$ $61.30$ $27.38$ $32.30$ $\mathbf{98.11}$
+$25\%$ 81.6081.60{81.60}81.60 79.6079.60{79.60}79.60 69.7069.70{69.70}69.70 78.5078.50{78.50}78.50 $65.40$ $\mathbf{50.00}$ $42.40$ $62.00$
+$40\%$ 61.6161.61{61.61}61.61 $61.30$ 31.3031.30{31.30}31.30 $61.60$ $\mathbf{72.60}$ $48.50$ $\mathbf{69.80}$ 83.3283.32{83.32}83.32
+MMD-AE $0\%$ $\mathbf{99.99}$ $\mathbf{99.29}$ 00.5300.53{00.53}00.53 $\mathbf{99.98}$ $\mathbf{55.47}$ $\mathbf{99.98}$ $\mathbf{99.69}$ $\mathbf{99.71}$
+$15\%$ 99.9899.98{99.98}99.98 02.5502.55{02.55}02.55 $\mathbf{22.55}$ 99.8399.83{99.83}99.83 12.1912.19{12.19}12.19 $41.34$ $98.77$ $98.67$
+$25\%$ $99.78$ 02.5202.52{02.52}02.52 $15.35$ $99.83$ $12.19$ $41.27$ $99.69$ $98.61$
+$40\%$ $99.96$ 01.8801.88{01.88}01.88 12.3012.30{12.30}12.30 $99.83$ $14.13$ 41.1241.12{41.12}41.12 56.2156.21{56.21}56.21 99.1499.14{99.14}99.14
+NSAE $0\%$ 99.9999.99{99.99}99.99 99.9899.98{99.98}99.98 00.0900.09{00.09}00.09 99.9999.99{99.99}99.99 $\mathbf{77.03}$ $\mathbf{99.90}$ $\mathbf{97.02}$ 99.5899.58{99.58}99.58
+$15\%$ 99.8099.80{99.80}99.80 $\mathbf{99.99}$ $04.82$ $34.92$ $36.04$ $00.00$ $05.16$ $\mathbf{99.80}$
+$25\%$ $\mathbf{99.99}$ $03.16$ 00.3200.32{00.32}00.32 $\mathbf{99.83}$ $59.36$ $15.90$ $00.35$ $99.69$
+$40\%$ 99.9899.98{99.98}99.98 34.3634.36{34.36}34.36 $\mathbf{53.66}$ $99.83$ 12.3612.36{12.36}12.36 00.0000.00{00.00}00.00 19.3319.33{19.33}19.33 $99.40$
+CORAL $0\%$ 59.1859.18{59.18}59.18 90.6190.61{90.61}90.61 30.4530.45{30.45}30.45 00.8000.80{00.80}00.80 $08.40$ $55.06$ $03.40$ $69.94$
+$15\%$ $61.53$ $12.64$ $\mathbf{31.79}$ $50.43$ $\mathbf{33.74}$ $41.31$ 50.3850.38{50.38}50.38 $67.54$
+$25\%$ $38.85$ $\mathbf{99.99}$ $0.00$ $0.01$ 22.9622.96{22.96}22.96 $82.21$ $0.00$ $95.01$
+$40\%$ $\mathbf{99.99}$ $38.85$ 00.0100.01{00.01}00.01 00.0000.00{00.00}00.00 $22.96$ $\mathbf{82.21}$ 00.0100.01{00.01}00.01 $\mathbf{99.19}$
+MTLS-RED $0\%$ $\mathbf{88.83}$ $\mathbf{86.08}$ $\mathbf{82.41}$ $\mathbf{96.12}$ $28.01$ $80.99$ $73.93$ $73.23$
+$15\%$ 79.0079.00{79.00}79.00 78.8578.85{78.85}78.85 78.7178.71{78.71}78.71 78.6878.68{78.68}78.68 $78.88$ $79.13$ $78.82$ $79.01$
+$25\%$ $83.90$ $80.70$ $80.70$ $83.80$ $76.50$ $\mathbf{81.96}$ $\mathbf{85.00}$ $77.10$
+$40\%$ 79.4679.46{79.46}79.46 79.3379.33{79.33}79.33 79.2079.20{79.20}79.20 79.1279.12{79.12}79.12 $\mathbf{78.99}$ 79.3379.33{79.33}79.33 79.2579.25{79.25}79.25 $\mathbf{79.64}$
+Table 1: Performance of the proposed method and baseline method with different percentages of cross-domain data added to the source domain during training.
+The results are on the CIC-CSE-IDS dataset.
+We highlight the best performance of each model across all the datasets.
 4.5 Ablation Study
-The ablation study is performed keeping the baseline models as reference. In order to study the effectiveness of our proposed model, we perform the following analysis
-  • We consider the multiple source and cross-domain reconstructions and use the encoded space to classify SOURCE, CROSS-DOMAIN, and OOD data.
-Our setup is different from [1] where the authors adopt a two-stage approach where the model is trained first, and then the trained encoder is used for classification purposes.
-On the other hand, our strategy is to jointly train all the components together so that the joint optimization with the de-correlation penalty can improve the generalization performance.
-The joint objective promotes discriminative power while simultaneously reducing the dependency on spurious dataset-specific correlation.
-  • In our experiments we increase the cross-domain data availability gradually from 5%-35% in the training data in order to understand the impact of decorrelation penalty in achieving domain generalization while reconstructing data from multiple domains.
-In Table 1, we observe that when the cross-domain attack classes constitute a small percentage $10\%$ of the source domain data, most of the baseline models overfit the source domain classes and do not perform well on the cross-domain and test OOD classes.
-The OOD classes are never involved in the training process.
-Whereas, our proposed model gives an overall good performance on all the source, cross-domain, and OOD classes except the RARE class.
-In Table 2, we report the results of the experiments where we gradually increase the cross-domain data to $35\%$.
-In this case, with our proposed model, the best accuracy generalizes well to all the source, cross-domain, and OOD classes although there is a decrease in performance with some of the source and cross-domain classes but better performance can be seen with the OOD classes especially.
-In a nutshell, we want to point out that by adjusting the kernel bandwidth and the sample size, we can get a reliable estimate of the mutual information penalty between prior and latent space which can help achieve better generalization to completely unseen classes that were never used in the training process.
+We perform different kinds of ablation studies to validate the performance of our model and also to compare our proposed model with standard baselines.
+In our experiments, we consider multiple source and cross-domain reconstructions and use the encoded space to classify source, cross-domain, and OOD data.
+Our setup is different from MTAE [1] where the authors adopt a two-stage approach where the model is trained first, and then the trained encoder is used for classification purposes.
+On the other hand, our strategy is to jointly optimize all the components together to improve the generalization performance.
+The joint objective promotes discriminative power while simultaneously reducing the dependency on spurious data-specific correlation.
+We try to design a principled approach to how cross-domain data can be combined with particular source domain data in order to improve the generalization to out-of-distribution data samples.
+We vary the cross-domain data in various proportions 5%-50% of the source data in the training data in order to understand the impact of the de-correlation penalty under different scenarios.
+In Table 1,2, we observe that as we gradually increase the percentage of the cross-domain classes in the training data, the performance of most of the baseline models deteriorates on the IN distribution.
+Whereas, with a small percentage of cross-domain data initially, most of the models overfit to the IN distribution datasets.
+The proposed model, on the other hand, gives an overall good performance on all the source, cross-domain, and OOD classes as we vary the amount of cross-domain data.
+It is to be noted that the OOD-domain data are never observed during any stage of the training process.
+In Table 1, we observe an interesting phenomenon where we gradually increase the cross-domain data to $30\%-40\%$, by adjusting the kernel bandwidth and the hyperparameters $\beta$ and $\lambda$, we can train a model that shows an overall generalization performance across all training and test datasets.
+The base models mostly overfit to the training data or a particular cross-domain data.
+In Table 2, we find that when we increase the proportion of cross-domain data during the training, the classification results are significantly good across the OOD datasets like WEB and SPOOFING attacks.
+Although there is a decrease in performance in some of the source domains, an overall improvement in performance across all classes can be seen.
+Determining the value of the kernel bandwidth parameter in kernel density estimation is important as it leads to over-smoothing or under-smoothing.
+In Figure 2, we plot the latent space of a simple multi-task encoder-decoder model without any regularization and our proposed model with the MI penalty.
+We see improved clustering with many of the source, cross, and target domain classes when the regularization is applied.
+Figure 2: Fig (f-j),(p-t) T-SNE projection of the latent space of DMTAE (without MI regularization) and (a-e),(k-o) is our proposed MTL-RED (with MI regularization).
+Again, (a-j) the attacks are GOLDEN-EYE, SOLARIS, BOTNET, HOIC, RARE.
+(k-t) the attacks are DOS, DDOS, MQTT, RECON, and SPOOFING.
+We see the difference distinctively in SOLARIS(b,g), RARE(e,j), DOS(k,p), DDOS(l,q) and MQTT(m,r).
+SOURCE DOMAIN CROSS DOMAIN OOD DOMAIN
+Model Percentage DDOS DOS RECON SPOOF MQTT MIRAI WEB BENIGN
+DMTAE $0\%$ $99.96$ 99.9999.99{99.99}99.99 54.0754.07{54.07}54.07 46.9846.98{46.98}46.98 $78.43$ $\mathbf{99.97}$ $30.55$ $97.53$
+$15\%$ $99.99$ 99.9999.99{99.99}99.99 98.6698.66{98.66}98.66 71.8471.84{71.84}71.84 $78.43$ $79.38$ $\mathbf{30.55}$ $97.53$
+$25\%$ 99.9999.99{99.99}99.99 99.9999.99{99.99}99.99 $\mathbf{99.99}$ $\mathbf{75.55}$ $\mathbf{99.93}$ $80.07$ $21.38$ $98.63$
+$40\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ 98.53498.534{98.534}98.534 $74.43$ $89.42$ $80.83$ $29.02$ $97.78$
+MMD-AE $0\%$ $99.96$ $99.96$ $49.95$ $41.32$ $\mathbf{95.87}$ $84.01$ $21.15$ $\mathbf{98.56}$
+$15\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ $98.47$ $70.54$ $90.54$ $\mathbf{84.01}$ $42.53$ $96.49$
+$25\%$ $99.99$ $99.99$ $99.11$ $73.60$ $89.42$ $77.60$ $58.04$ $93.48$
+$40\%$ $99.61$ $99.46$ $\mathbf{99.30}$ $\mathbf{78.23}$ $94.84$ $72.08$ $\mathbf{67.21}$ $92.41$
+NSAE $0\%$ $\mathbf{99.99}$ $99.99$ $46.90$ $32.42$ $69.11$ $\mathbf{99.90}$ $\mathbf{56.05}$ $94.68$
+$15\%$ $99.99$ $\mathbf{99.99}$ $97.77$ $69.11$ $\mathbf{99.90}$ $99.87$ $36.66$ $\mathbf{97.71}$
+$25\%$ $99.99$ $99.99$ $\mathbf{98.70}$ $\mathbf{71.29}$ $99.71$ $99.90$ $43.83$ $96.11$
+$40\%$ $99.99$ $99.99$ $98.62$ $70.64$ $99.89$ $99.90$ $49.70$ $95.78$
+CORAL $0\%$ $\mathbf{99.99}$ $99.99$ $98.53$ $74.43$ $89.42$ $\mathbf{81.13}$ $30.53$ $98.73$
+$15\%$ $99.99$ $\mathbf{99.99}$ $98.53$ $74.43$ $89.42$ $81.13$ $\mathbf{42.30}$ $98.73$
+$25\%$ $99.99$ $99.99$ $\mathbf{98.86}$ $75.55$ $\mathbf{99.90}$ $80.08$ $42.03$ $\mathbf{98.73}$
+$40\%$ $99.99$ $99.99$ $99.11$ $\mathbf{77.21}$ $99.57$ $80.51$ $42.30$ $98.73$
+MTLS-RED $0\%$ $99.99$ $99.99$ $98.41$ 78.9278.92{78.92}78.92 $78.21$ 76.0476.04{76.04}76.04 $73.91$ $87.67$
+$15\%$ $99.99$ $99.99$ $98.41$ 78.9278.92{78.92}78.92 $78.21$ 99.7599.75{99.75}99.75 $68.03$ $91.83$
+$25\%$ $99.99$ $99.99$ $98.41$ $78.21$ $53.19$ 99.8799.87{99.87}99.87 $73.67$ $93.05$
+$40\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ $\mathbf{98.98}$ $\mathbf{81.17}$ $\mathbf{95.87}$ $\mathbf{99.88}$ $\mathbf{75.91}$ $\mathbf{91.2}$
+Table 2: Performance of the proposed method and baseline method with different percentages of cross-domain data added to the source domain during training.
+The results are on the combined features of CIC-IOT and CIC-IoMT datasets.
 5 Conclusion
-In this paper, we attempt to study an important problem of domain generalization where a model trained on several sources and cross-domains can generalize better on OOD classes(’zero-day’ attacks in the case of cyber-security).
-We showcase that learning feature representation from multiple diverse domains in the latent space with proper regularization can achieve better OOD generalization.
-We formulate a joint objective that tries to classify data in a multi-domain latent space while simultaneously enforcing a de-correlation penalty in the form of mutual information between input and latent space to create a domain-invariant latent space.
+In this paper, we attempt to study an important problem of domain generalization where a model trained on several sources and cross-domains can generalize better on OOD classes.
+We showcase that learning feature representation from multiple diverse domains in the latent space with proper regularization designed to remove spurious correlation can achieve better OOD generalization.
+We formulate a joint objective that tries to classify data in a multi-domain latent space while simultaneously enforcing a de-correlation penalty in the form of mutual information between input and latent space in order to create a domain-invariant latent space.
 Through detailed experiments, we showcase the performance of the model on benchmark cybersecurity datasets with the help of standard classification metrics.
-In the future, we would like to study domain generalization in time series and other sequential data using latent space regularization.
+Since the kernel bandwidth is sensitive to the mutual information estimation, it would be interesting to learn this parameter during the training process.
+In the future, we would like to study domain generalization in time series and other sequential data with latent space techniques.
+Appendix 0.A Intuition behind the Mutual Information Penalty
+We consider a neural network encoder $E$, encoding input $X$ to latent space $Z$.
+Here, our goal is to remove the dependency of the encoder of the spurious correlation that is inherited from the high dimensional input space $X$ to the latent space $Z$.
+To achieve this, we augment the multi-task cross-entropy objective with mutual information between the original input space $X$ and latent space $Z$ which aims to create a disentangled or invariant latent space.
+$\mathcal{L}\left(\phi,\theta;X,Z\right)=\mathcal{L}_{ce}+\beta MI\left(X,Z\right)$ (7)
+where, $\mathcal{L}_{ce}$ is the generic classification loss.
+$MI(X,Z)$ is the mutual information loss between the encoding $Z$ and input space $X$.
+We aim to penalize large spurious information in the encoding space $Z$ and the penalty strength is determined by the parameter $\beta$.
+We consider the scenario where we do not know the specific features in the domains that have spurious correlation beforehand, therefore we directly minimize the MI content in the kernel space of the input and the latent representations which indirectly will aim to remove the spurious correlation in the high dimensional prior space and create a disentangled latent space.
+The mutual information(MI) content between the latent space and the input space can be expressed in the form of their joint and marginal distributions.
+By enforcing decorrelation, we minimize the MI divergence between joint and the product of the marginal distributions.
+Since MI is intractable to compute in general, the lower bound on the MI [38] can be given by:
+$\begin{split}MI\left(E_{\phi}\left(X\right),Z\right)&=\mathbf{E}_{P_{X,Z}}\left[D_{\theta}\left({E_{\phi}\left(X\right),Z}\right)\right]-\log\mathbf{E}_{P_{X}\bigotimes P_{Z}}\left[e^{D_{\theta}\left({E_{\phi}\left(X\right)},Z\right)}\right]\\ &\leq MI\left(E_{\phi}\left(X\right),Z\right)\end{split}$
+The final objective can be written as,
+$\min_{\theta,\phi}\mathbf{E}\left[-\log p_{\theta}\left(Y|X\right)\right]+\beta MI\left({E}_{\phi}\left(X\right),Z\right)$ (8)
+The classifier or the encoder tries to enforce de-correlation by minimizing the $MI\left(X,Z\right)$ in the kernel space of the input $\left(\mathcal{K}_{x}\right)$ and the latent representations $\left(\mathcal{K}_{z}\right)$ simultaneously with the cross-entropy loss.
+$\beta$ controls the trade-off between the de-correlation and the classification loss.
+The numerical stable objective is the sum of these two loss functions.
+The first part promotes discrimination power while the second part reduces the classification dependence on the spurious features in $X$.
+Appendix 0.B Differentiability of the Mutual Information Loss
+Considering Renyi’s second-order entropy where $\alpha=2$, the entropy or the uncertainty of the input space can be estimated as
+$\hat{H_{2}}(\mathcal{K}_{X})=-\log_{2}(tr(\mathcal{K}_{X}^{2}))=-\log_{2}\left(\sum_{i=1}^{N}\lambda_{i}(\mathcal{K}_{X})^{2}\right),$ (9)
+where, $\lambda_{i}(\mathcal{K}_{X})$ denotes the $i^{th}$ eigenvalue of $\mathcal{K}_{X}$.
+We used this measure to estimate the entropy of the original input data space.
+Similarly, the entropy of the latent space can be estimated as
+$\hat{H_{2}}(\mathcal{K}_{Z})=-\log_{2}(tr(\mathcal{K}_{Z}^{2}))=-\log_{2}\left(\sum_{i=1}^{N}\lambda_{i}(\mathcal{K}_{Z})^{2}\right),$ (10)
+. Now, the joint entropy between input and latent space can be measured as,
+$\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)=H_{2}\left(\frac{\mathcal{K}_{X}\circ\mathcal{K}_{Z}}{tr\left(\mathcal{K}_{X}\circ\mathcal{K}_{Z}\right)}\right),$ (11)
+where $\circ$ denotes the Hadamard product of the matrices. Taking the analytical gradients, we can write:
+$\frac{\partial H_{2}\left(\mathcal{K}_{X}\right)}{\partial X}=-2\frac{\mathcal{K}_{X}}{tr\left(\mathcal{K}_{X}^{2}\right)}$ (12)
+Similarly, the joint entropy between prior and latent space can be estimated as
+$\frac{\partial H_{2}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)}{\partial X}=-2\left[\frac{\left(\mathcal{K}_{X}\circ\mathcal{K}_{Z}\right)\circ\mathcal{K}_{Z}}{tr\left(\mathcal{K}_{X}\circ\mathcal{K}_{Z}\right)^{2}}-\frac{MI\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)\circ\mathcal{K}_{Z}}{tr\left(\mathcal{K}_{X}\circ\mathcal{K}_{Z}\right)}\right]$ (13)
+where, $MI\left(X,Z\right)$ is the mutual entropy between the prior and the latent space.
+Now, the conditional entropy between latent and prior space can be estimated as:
+$\frac{\partial MI_{2}\left(\mathcal{K}_{X};\mathcal{K}_{Z}\right)}{\partial X}=\frac{\partial H_{2}\left(\mathcal{K}_{X}\right)}{\partial X}+\frac{\partial H_{2}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)}{\partial X}$ (14)
+Appendix 0.C Selecting the cross-domains, source and OOD domains
+We consider a setting where the joint distribution of the data doesn’t have much overlap but the marginal distribution of the features do have an overlap.
+This scenario happens when the features exhibit different correlations between the source and cross-domain data distributions.
+For example, we see that DOS and DDOS datasets have very similar correlations between the features.
+However, the correlation between the features in the other attacks like MIRAI and WEB is different.
+Similarly, the GOLDEN and SOLARIS datasets have a certain kind of correlation between the features that differ from other attack datasets like INFIL and BOTNET.
+Our goal is to combine the datasets with different patterns of correlation structure among its features during the training phase in order to increase the versatility of the latent space.
+Figure 3: Fig (a-o) Correlation plots of various cyber-security datasets.
+(a)DOS-IOT (b) DDOS IOT (c) WEB (d) MIRAI (e) BENIGN IOT (f) DDOS IOMT (g) DDOS IOMT (h) RECON IOMT (I) BENIGN IOMT (j) BENIGN CIC IDS (k) GOLDEN CIC IDS (l) HOIC (m) INFIL CIC IDS (n) RARE CIC IDS (0) SOLARIS CIC IDS.